### v1

# Properties of current sheets in two-dimensional tearing-mediated magnetohydrodynamic turbulence

Abstract
It is well known that the nonlinear evolution of magnetohydrodynamic (MHD) turbulence generates intermittent current sheets. In the solar wind turbulence, current sheets are frequently observed and they are believed to be an important pathway for the turbulence energy to dissipate and heat the plasma. In this study, we perform a comprehensive analysis of current sheets in a high-resolution two-dimensional simulation of balanced, incompressible MHD turbulence. The simulation parameters are selected such that tearing mode instability is triggered and plasmoids are generated throughout the simulation domain. We develop an automated method to identify current sheets and accurately quantify their key parameters including thickness ($a$), length ($L$), and Lundquist number ($S$). Before the triggering of tearing instability, the current sheet lengths are mostly comparable to the energy injection scale. After the tearing mode onsets, smaller current sheets with lower Lundquist numbers are generated. We find that the aspect ratio ($a/L$) of the current sheets scales approximately as $S^{-1/2}$, i.e. the Sweet-Parker scaling.While a power-law scaling between $L$ and $a$ is observed, no clear correlation is found between the upstream magnetic field strength and thickness $a$. Finally, although the turbulence energy shows anisotropy between the directions parallel and perpendicular to the local magnetic field increment, we do not observe a direct correspondence between the shape of the current sheets and that of the turbulence “eddies.” These results suggest that one needs to be cautious when applying the scale-dependent dynamic alignment model to the analysis of current sheets in MHD turbulence.
software: Matplotlib (J. D. Hunter, 2007), NumPy (C. R. Harris et al., 2020), Scikit-Learn (F. Pedregosa et al., 2011), LAPS (C. Shi et al., 2024)
1 Introduction
It has long been observed that the solar wind is highly turbulent, containing fluctuations on a vast span of scales (R. Bruno & V. Carbone, 2013). The solar wind turbulence is believed to be one of the major contributors to the heating and acceleration of solar wind and thus has been extensively investigated for decades (e.g. E. Marsch & C.-Y. Tu, 1990; B. Goldstein et al., 1995; R. Grappin & M. Velli, 1996; W. Matthaeus et al., 2005; C. H. Chen et al., 2012; O. Alexandrova et al., 2013; C. Shi et al., 2021; N. Sioulas et al., 2023).
One of the key questions regarding solar wind turbulence is how its energy is dissipated. The widely-accepted picture is that, as the fluctuation energy cascades from large, MHD scales toward small, kinetic scales, various types of wave-particle interactions take effect, transferring energy from the turbulence to the particles (e.g. S. S. Cerri et al., 2017; J. He et al., 2019; M. M. Martinović et al., 2020; T. A. Bowen et al., 2022). In addition to these wave-particle interaction processes, intermittency naturally arises in turbulence systems, including neutral fluid (e.g. U. Frisch & G. Parisi, 1980) and plasma (e.g. L. Sorriso-Valvo et al., 2001). Intermittency refers to the sparse, irregular and strong fluctuations on top of the periodic and smooth part of the turbulence, and it breaks down the self-similarity of the turbulence. In MHD turbulence, intermittency mainly appears in the form of current sheets and vortices (M. Wan et al., 2016), and it is well-known that current sheets are usually accompanied by strong dissipation via magnetic reconnection (K. Osman et al., 2014). Therefore, intermittency may play a significant role in dissipation of MHD turbulence. In the solar wind, intermittency has been extensively observed (R. Bruno et al., 2001; W. H. Matthaeus et al., 2015; Z. Pei et al., 2016; H. Wu et al., 2023), and evidence shows that intermittency is co-located with regions of higher proton and electron temperatures in the solar wind (K. Osman et al., 2012; N. Sioulas et al., 2022b, a).
The question then comes to how do we understand the generation of intermittency in MHD turbulence. In incompressible-MHD, phenomenological models were developed based on the idea of scale-dependent dynamic alignment (SDDA) (S. Boldyrev, 2005; J. Podesta et al., 2009; B. D. Chandran et al., 2015; A. Mallet et al., 2016; N. Sioulas et al., 2024a). In this scenario, the small-scale turbulence “eddies” are assumed to be anisotropic three-dimensionally. Let us write the eddy size along the background magnetic field $\bm{B_{0}}$ as $l$, the eddy size along the polarization direction of the fluctuation $\bm{\delta b}$ in the plane perpendicular to $\bm{B_{0}}$ as $\xi$, and the eddy size perpendicular to both $\bm{B_{0}}$ and $\bm{\delta b}$ as $\lambda$. In the perpendicular plane, it was assumed that the eddies are extended along the polarization direction, i.e. $\xi>\lambda$. A set of scaling relations between these lengths were derived, which were then coupled with the tearing instability theory with the assumption that the ratio $\xi/\lambda$ corresponds to the aspect ratio of the small-scale current sheets (S. Boldyrev, 2006; S. Boldyrev & N. F. Loureiro, 2017; N. F. Loureiro & S. Boldyrev, 2017; A. Mallet et al., 2017). As the scale decreases, the aspect ratio may reach a certain critical value, after which the turbulence is dominated by the recursive collapse of current sheets, i.e. the so-called “reconnection-mediated” regime, associated with a steepened power spectrum (C. Dong et al., 2022).
Although both observations (e.g. J. Podesta et al., 2009; N. Sioulas et al., 2024a) and numerical simulations (e.g. J. Mason et al., 2006; J. Walker et al., 2018) have shown evidence of SDDA in plasma turbulence, there is a potential caveat in applying the SDDA theory to the reconnection in turbulence. That is, the turbulence-generated current sheets are intermittent and thus are sparse in space instead of being space-filling like the turbulent eddies. Hence, whether the anisotropic eddies directly correspond to the shape of intermittent current sheets is questionable. In this paper, we conduct and analyze a high-resolution 2D incompressible MHD simulation of balanced turbulence. The main objective is to compare the scale-dependent dynamic alignment with the properties of intermittent current sheets. We emphasize that, the classic SDDA theories are not strictly applicable to 2D turbulence due to the lack of propagation effect in 2D ($l\rightarrow+\infty$). Nonetheless, we are still able to compute the aspect ratio of the 2D eddies $\xi/\lambda$ and verify whether they are related to the aspect ratio of the current sheets.
The paper is organized as follows: In Section 2, we introduce the numerical methods and describe the simulation setup. In Section 3, we present the results, including the global evolution of turbulence properties (Section 3.1), the method to identify and analyze the intermittent current sheets (Section 3.2), a case study of an evolving current sheet (Section 3.3), and the statistics of the current sheet properties (Section 3.4). In Section 4, we discuss the results with a focus on comparing SDDA with current sheet properties. In Section 5, we summarize this work.
2 Simulation setup
We use the UCLA-Pseudo-Spectral (LAPS) code to conduct the simulation. LAPS is a Fourier-transform based pseudo-spectral MHD code (C. Shi et al., 2024). It has compressible and incompressible versions implemented in both 2D and 3D. Here, we use the incompressible version which allows us to compare the results with most previous theoretical models that were developed based on the assumption of incompressibility. We carry out the simulation in 2D so that high resolution can be adopted, which is essential to the development of tearing instability.
The simulation domain (in xyx-y plane) has the size $L_{x}=L_{y}=1.0$ and numbers of grid points $N_{x}=N_{y}=8192$. The density is uniform $\rho\equiv 1$ and we set an out-of-plane magnetic field $B_{z}\equiv 1$. We note that in incompressible 2D MHD, $B_{z}$ does not affect the evolution and it only serves as a normalization unit in this case. We initialize the simulation by adding Alfvén-type fluctuations on all wave modes satisfying $k_{min}\leq\left|k\right|\leq k_{max}$, where $k_{min}=8$ and $k_{max}=16$, thus the initial power spectrum is isotropic. Here $k=1/\lambda$ with $\lambda$ being the wavelength.
Figure 1: Time evolution of (a) $\sigma_{c}$ (blue) and $\sigma_{r}$ (orange); (b) kinetic energy (blue), magnetic energy (orange), and energies of $\bm{z^{+}}$ (green) and $\bm{z^{-}}$ (red); (c) averaged $J^{2}$ (blue), $\omega^{2}$ (orange), and $2\bm{J}\cdot\bm{\omega}$ (green).
Figure 2: Power spectra of (a) magnetic field, (b) velocity, (c) $\bm{z^{+}}$, and (d) $\bm{z^{-}}$ calculated along $x$. In each panel, different solid curves correspond to spectra at different time moments. Blue dotted-dashed lines are linear-fitting of the spectra at $t=0.6$ using $\frac{1}{128}\leq k\leq\frac{1}{16}$. The black dashed lines show $\propto k^{-5/3}$ for reference. The yellow shade in each panel marks $\frac{1}{16}\leq k\leq\frac{1}{8}$ that correspond to the initial fluctuations.
For each $\bm{k}$, the perturbation satisfies $\bm{b_{k}}\perp\bm{k}$, and we add the two Alfvén modes, i.e. $\bm{v_{k}}=\pm\bm{b_{k}}$, with the same amplitude but random phases. Consequently, the normalized cross helicity
$\sigma_{c}=\frac{E_{+}-E_{-}}{E_{+}+E_{-}}$
and normalized residual energy
$\sigma_{r}=\frac{E_{v}-E_{b}}{E_{v}+E_{b}}$
are both approximately zero. Here $E_{v}=\int v^{2}$, $E_{b}=\int b^{2}$, and $E_{\pm}=\int{z^{\pm}}^{2}$ measure the total energies in velocity, magnetic field, and the two Elsässer variables $\bm{z^{\pm}}=\bm{v}\mp\bm{b}$. The root-mean-square amplitude of the initial fluctuation is $RMS(\bm{b})=0.173$. Explicit resistivity $\eta$ and viscosity $\nu$ are implemented with $\eta=\nu=5\times 10^{-7}$. Thus the (magnetic) Reynolds number is estimated to be $S_{0}=\lambda\delta b/\eta\approx 4\times 10^{4}$. We conducted tests with different amplitudes of dissipation and the values $\eta=\nu=5\times 10^{-7}$ are chosen such that the dissipation is strong enough to effectively suppress the numerical error induced by Gibbs phenomenon while not too strong so that a significant number of tearing-unstable current sheets are generated (M. Wan et al., 2013). The eddy-turnover time $\tau_{nl}=\lambda/z_{\lambda}\approx 0.7$. Thus, we run the simulation to $t=1.0$, roughly 1.4 eddy-turnover times.
Figure 3: Evolution of $J_{z}$ in the subdomain $x\in[0.2,0.7]$, $y\in[0,0.5]$. In Panel (c), the green box marks the current sheet analyzed in detail in Section 3.3 and Figure 6.
3 Results
3.1 Evolution of turbulence properties
In Figure 1, we show the time evolution of various quantities. Panel (a) shows that $\sigma_{c}$ (blue) remains roughly zero throughout the simulation, while $\sigma_{r}$ (orange) decreases rapidly at the early stage ($t\leq 0.15$), then increases a little bit and finally starts to decay slowly. The initial rapid decrease of $\sigma_{r}$ was observed in 3D simulations with similar setup (C. Shi et al., 2025). In Panel (b), we show the evolution of $E_{v}$ (blue), $E_{b}$ (orange), $E_{+}$ (green), and $E_{-}$ (red) respectively. $E_{+}$ and $E_{-}$ decay slowly due to the numerical dissipation, and so for $E_{v}$ and $E_{b}$ after $t\approx 0.3$. The early stage ($t\leq 0.15$) is featured by a rapid increase in $E_{b}$ and decrease in $E_{v}$, leading to the decrease in $\sigma_{r}$. We mark the peak of $E_{b}$ (and the valley of $E_{v}$) with the vertical dotted line. Clearly, there is a fast transfer of kinetic energy to magnetic energy at the early stage, much earlier than one eddy-turnover time. As will be shown in Figure 3, thin current sheets are formed concurrently. After $t=0.15$, reconnection becomes significant in these current sheets, transferring magnetic energy back to the kinetic energy, manifested by the slight increase in $E_{v}$. We note that, the negative residual energy is concentrated near the eddy boundaries, in regions surrounding the thin current sheets, rather than inside the current sheets where plasma jets are generated and thus featured by positive residual energy.
Panel (c) shows evolution of energy dissipation rates, i.e. the magnetic energy dissipation rate $\langle|\bm{J}|^{2}\rangle$ (blue) with $\bm{J}=\nabla\times\bm{b}$ being the current density, the kinetic energy dissipation rate $\langle|\bm{\omega}|^{2}\rangle$ (orange) with $\bm{\omega}=\nabla\times\bm{v}$ being the vorticity, and the cross term $2\langle\bm{J}\cdot\bm{\omega}\rangle$ (green). In this simulation, $\bm{J}=J_{z}\hat{e}_{z}$ and $\bm{\omega}=\omega_{z}\hat{e}_{z}$. We note that the dissipation rates for $\bm{z^{\pm}}$ are $\langle|\bm{J}|^{2}+|\bm{\omega}|^{2}\mp 2\bm{J}\cdot\bm{\omega}\rangle$ respectively. Here $\langle\rangle$ represents average over the whole simulation domain. The vertical dashed line marks $t=0.6$, corresponding to the peaks of $\langle\bm{J}^{2}\rangle$ and $\langle\bm{\omega}^{2}\rangle$. This is the time moment when the turbulence is fully developed. After this moment, the turbulence starts to decay. Note that 0.6 is close to one eddy turnover time $\tau_{nl}\approx 0.7$.
Figure 4: Scale-dependent Kurtosis of magnetic field at different time moments. The yellow shade marks the wavelength-range of the initial fluctuations.
Figure 5: (a) Same as Panel (c) of Figure 3 with identified current sheets marked by yellow ($J_{z}>0$) and green ($J_{z}<0$). (b) Probability distribution function (PDF) of $J_{z}$ at $t=0.3$. Blue bars are all the data points and orange bars are data points with identified current sheets removed. (c) Time evolution of filling factor (blue), i.e. area of current sheets divided by the total area of the domain, and number of current sheets (orange). (d) $\int J_{z}^{2}$ inside the current sheets over $\int J_{z}^{2}$ throughout the simulation domain.
In Figure 2, we show the power spectra for different quantities at different time moments. We fit the spectra at $t=0.6$, when the turbulence is fully developed and the spectra are stably established over $\frac{1}{128}\leq k\leq\frac{1}{16}$, which is shown by the blue dotted-dashed lines. The fitted slopes are written in the legends. The magnetic field spectral slope is close to 5/3-5/3 and the velocity spectrum is shallower. This has been confirmed by a large number of numerical simulations (e.g. W.-C. Müller & R. Grappin, 2005; S. Boldyrev et al., 2011; C. Shi et al., 2025) and observations (e.g. J. Podesta et al., 2007; C. Chen et al., 2013; C. Shi et al., 2021). The slopes for $\bm{z^{\pm}}$ are similar and take values between 3/2-3/2 and 5/3-5/3. Figure 2 confirms that the formation of thin current sheets ($t=0.15$, Figure 1) is well before the turbulence is fully developed, since the power spectra are still evolving at $t=0.15$.
In Figure 3, we show four snapshots of $J_{z}$ in the subdomain $x\in[0.2,0.7],\,y\in[0,0.5]$. Panel (a) shows the initial condition, where $J_{z}$ is smooth (note that the color bar is in log-scale). Panel (b) shows $t=0.15$, corresponding to the vertical dotted line in Figure 1. At this moment, a number of thin current sheets are generated, but fast reconnection has not onset yet. Panel (c) shows $t=0.3$, when more current sheets are generated and the current sheets become thinner than $t=0.15$. Tearing instability has been triggered in some of the current sheets, indicated by the generation of plasmoids. One tearing-unstable current sheet is marked by the green box and this current sheet will be analyzed in detail in Section 3.3. Panel (d) shows $t=0.6$, corresponding to the vertical dashed line in Figure 1, when the turbulence is fully developed. Much more plasmoids are generated at this time compared with $t=0.3$. Figure 3 shows that the generated current sheets are located near the boundaries of eddies whose sizes are close to to the wavelengths of the initial fluctuations.
In Figure 4, we show the scale dependent Kurtosis of the magnetic field, defined by
$\mathrm{Kurtosis}(\bm{b},l)=\frac{\langle\left|\delta\bm{b}(l)\right|^{4}\rangle}{\langle\left|\delta\bm{b}(l)\right|^{2}\rangle^{2}},$ (1)
at different time moments. Here $\delta\bm{b}=\bm{b}(\bm{x}+\bm{l})-\bm{b}(\bm{l})$. In general, the smaller scales correspond to larger values of Kurtosis. Interestingly, it is observed that the Kurtosis reaches maximum at $t=0.15$ and then gradually decay. Combined with Figure 1, it indicates that the increase of Kurtosis is associated with the current sheet formation process, and the later reconnection decreases the Kurtosis.
3.2 Identification of current sheets
To quantify the properties of the current sheets, it is necessary to accurately identify the current sheets. We develop an algorithm similar to that described in (V. Zhdankin et al., 2013), and we briefly describe our algorithm here. We start from finding the grid point corresponding to $\max(\left|J_{z}\right|)$ throughout the whole domain, excluding regions already marked as current sheets. Then we run a recursive, depth-first search algorithm to check the four neighboring grid points. The algorithm returns when all the neighboring points have $\left|J_{z}\right|\leq 0.1\max(\left|J_{z}\right|)$, and we mark the continuous set of data points as one single current sheet. This current-sheet search is repeated until all the current sheets with peak-$\left|J_{z}\right|$ larger than $0.1\max_{global}(\left|J_{z}\right|)$ are identified, where $\max_{global}(\left|J_{z}\right|)$ is the maximum $\left|J_{z}\right|$ in the whole simulation domain. In Figure 5(a), we show the same plot as Figure 3(c) and mark the identified current sheets in yellow ($J_{z}>0$) and green ($J_{z}<0$). It shows that the algorithm works well, and most of the identified current sheets exist between the interacting large-scale “eddies”, i.e. near the eddy boundaries. However, we note that, as the turbulence develops, there are many regions with strong current density but are “broken” instead of being coherent structures. Thus, we further refine the current sheet lists by discarding regions with less than 20 data points.
In Panel (b) of Figure 5, we show the probability distribution function (PDF) of $J_{z}$ at $t=0.3$. Here blue bars are all the data points in the whole simulation domain and orange bars are points after removing the identified current sheets. The PDF without current sheets is naturally thinner than that with current sheets included. Panel (c) displays the filling factor (blue) and the number (orange) of current sheets. The filling factor is defined as the total area of the identified current sheets divided by the area of the simulation domain. Panel (d) shows the normalized contribution of the identified current sheets to the total magnetic energy dissipation, i.e. ratio between the integrated $J_{z}^{2}$ in the current sheets and the integrated $J_{z}^{2}$ throughout the whole simulation domain. Panels (c) and (d) show that, the current sheets, with a total area less than 10% of the simulation domain, contribute roughly 50% of the dissipation of magnetic energy.
Figure 6: Time evolution of a single current sheet within the domain marked by the green box in Figure 3. In each panel, top plot shows $J_{z}$ and bottom plot shows $\omega_{z}$. In panels (a) & (b), the green shade is the automatically identified current sheet region using the algorithm described in Section 3.2. The white and red lines correspond to the segments of the current sheet given by the recursive PCA algorithm described in Section 3.3.
3.3 Case study of a current sheet
We aim to accurately quantify the geometry of these current sheets, including their lengths and thicknesses. Therefore, a principal component analysis (PCA) based algorithm is developed, with the aid of PCA functions implemented in the Python package Scikit-Learn (F. Pedregosa et al., 2011).
For a set of spatial coordinates $(x_{i},y_{i})$ ($i=0,\cdots N-1$), PCA determines two orthogonal unit vectors, along which the set of coordinates has maximum and minimum variances respectively. We refer to the axes along the two unit vectors as major (maximum variance) and minor (minimum variance) axis. For a straight current sheet, the major and minor axes roughly correspond to its length and thickness. However, as can be seen from Figure 5(a), most of the current sheets are curved, hence increasing the error in quantifying the current sheet length. To reduce this error, we apply PCA recursively for each current sheet. That is, after applying PCA, we check the ratio between the maximum variance and minimum variance. If this ratio is larger than 3, it means the data points are still quite spread in a long and thin region, and we further divide the dataset into two subsets, i.e. left and right of the center of mass $(\langle x_{i}\rangle,\langle y_{i}\rangle)$ along the major axis, and apply PCA to the two subsets separately until the ratio between maximum and minimum variances gets smaller than 3. With this recursive-PCA algorithm, we are able to break a long and curved current sheet into multiple segments along the major axis. For each segment, we define the length as the extent of the data points along the major axis, and the length of the current sheet is the sum-up of the lengths of all the segments.
Figure 7: (a) $B_{l}$ as a function of distance along the minor axis at $t=0.25$. Different curves correspond to different segments (see the text) in the recursive PCA analysis. Red circles mark the maximum and minimum of each $B_{l}$ profile. (b) $V_{l}$ as a function of distance along the major axis, at the center of the current sheet, at $t=0.25$. (c) Time evolution of the length (blue circle) and half-width (orange square) of the current sheet. (d) Time evolution of the upstream magnetic field strength $B_{up}$ and the outflow speed $V_{out}$. (e) Time evolution of the Lundquist numbers $S_{L}=LB_{up}/\eta$ (blue circle) and $S_{a}=aB_{up}/\eta$ (orange square).
In Figure 6, we show evolution of the current sheet marked by the green box in Figure 3(c). At each time moment, we plot $J_{z}$ on the top and $\omega_{z}$ on the bottom. In panels (a) & (b), the identified current sheet region is marked by the green shade. The short white and red lines correspond to the segments given by the recursive PCA algorithm. Clearly, from $t=0.2$ to $t=0.25$, the current sheet is stretched and thinned. The vorticity shows a quadrupole structure, implying bidirectional jets, which is evidence of ongoing reconnection. At $t=0.25$, signature of shear-driven instability, i.e. Kelvin-Helmholtz instability, is observed at the tips of the current sheet. At $t=0.3$, tearing instability is triggered, generating multiple magnetic islands. We estimate the average length and half-width of the magnetic islands to be $\lambda\approx 5\times 10^{-3}$ and $w\approx 5\times 10^{-4}$ respectively. Later, at $t=0.35$, the tearing instability enters nonlinear stage, featured by coalescing magnetic islands.
In Figure 7, we show the PCA result of the current sheet shown in Figure 6. In panel (a), we show $B_{l}$, i.e. magnetic field projected along the major axis, as a function of the coordinate along minor axis at $t=0.25$. Each curve corresponds to one segment in the PCA analysis. One can see that the $B_{l}$ profile is similar to Harris type but its amplitude decreases far away from the center of the current sheet. For each profile, we mark the maximum/minimum of $B_{l}$ with the red circles, and take the distance between the two red circles as the width (thickness) of the current sheet segment. We then average the widths of all the segments and define this average value as the width of the current sheet. In panel (b), we show the variation of $V_{l}$ along the major axis at the center of the current sheet at $t=0.25$. $V_{l}$ changes from 0.3-0.3 to $0.4$ as we move from the left to the right side of the current sheet. Panel (c) shows the time evolution of the length $L$ (blue circles) and half-width $a$ (orange squares) of the current sheet. For $t\leq 0.25$, the current sheet lengthens and thins until $L\approx 0.08$ and $a\approx 0.005$, after which it enters the nonlinear tearing stage. Panel (d) shows the time evolution of the upstream magnetic field $B_{up}$ (blue), defined as $\langle B_{l,max}-B_{l,min}\rangle/2$ where $B_{l,max}$ and $B_{l,min}$ correspond to the red circles and $\langle\rangle$ means average over all segments, and the outflow speed $V_{out}$, defined as $(V_{l,max}-V_{l,min})/2$ along the major axis as shown by panel (b). Before the nonlinear tearing stage, the two quantities are similar. As the tearing mode grows, the reconnection is no longer laminar, and $V_{out}$ becomes smaller than $B_{up}$. Last, in panel (e), we show the evolution of the estimated Lundquist numbers $S_{L}=LB_{up}/\eta$ and $S_{a}=aB_{up}/\eta$. In the linear-tearing stage ($t=0.25$), we have $S_{L}\approx 6.5\times 10^{4}$ and $S_{a}\approx 500$. Note that at this moment $a/L\approx 5\times 10^{-4}/0.08=0.00625\approx S_{L}^{-0.46}$, close to the Sweet-Parker model which predicts $a/L\sim S_{L}^{-1/2}$.
In summary, we conclude that this single current sheet undergoes the following stages: (1) Formation by thinning and lengthening accompanied by laminar reconnection. (2) Growth of linear tearing instability when the aspect ratio approaches the Sweet-Parker scaling $a/L\sim S_{L}^{-1/2}$. (3) Fast dissipation of the current sheet by nonlinear tearing instability. We note that, the Lundquist number $S_{L}=6.5\times 10^{4}$ of this current sheet is not significantly larger than the critical Lundquist number, approximately $O(10^{4})$, for the Sweet-Parker type current sheet to be tearing-unstable (e.g. D. Biskamp, 1986; N. Loureiro et al., 2007; C. Shi et al., 2018). In the nearly-collisionless space plasma with much larger Lundquist numbers, the growth of linear tearing mode may happen earlier, when the aspect ratio $a/L$ approaches $S_{L}^{-1/3}$ instead of $S_{L}^{-1/3}$ (F. Pucci & M. Velli, 2013; A. Tenerani et al., 2015, 2016).
3.4 Statistics of current sheets
Figure 8: Statistics of the current sheet properties. Top to bottom rows show $t=0.15$, $0.30$, and $0.60$ respectively. Left column (a) shows the aspect ratio $a/L$ as a function of Lundquist number $S_{L}$. Middle column (b) shows the upstream magnetic field $B_{up}$ as a function of half-width of the current sheet $a$. Right column (c) shows the current sheet length $L$ as a function of its half-width $a$. Each dot corresponds to one single current sheet and is color-coded with the peak current density inside the current sheet. In the right column, the yellow shades mark $1/16\leq L\leq 1/8$.
We apply the analysis described in Section 3.3 to all the identified current sheets and present the results in Figure 8. Top to bottom rows show time moments $t=0.15$, $0.3$, and $0.6$ respectively. Left column shows aspect ratio $a/L$ versus Lundquist number $S_{L}$, middle column shows upstream magnetic field $B_{up}$ versus current sheet thickness $a$, and right column shows current sheet length $L$ versus thickness $a$. Each dot corresponds to one identified current sheet, and the dot is color-coded by the peak current density in the current sheet. In the left column, we plot reference lines $a/L\propto S_{L}^{-1/2}$, i.e. the Sweet-Parker type current sheet model. We can see that for all the three moments, the data points show a negative correlation between $a/L$ and $S_{L}$, and they roughly obey $a/L\sim S_{L}^{-1/2}$. In the middle and right columns, we plot reference lines $B_{up}\propto a^{1/4}$ and $L\propto a^{3/4}$, which are predictions by the SDDA theory (S. Boldyrev, 2006). There is a strong positive correlation between $L$ and $a$, and they may roughly follow the power-law relation $L\propto a^{3/4}$. In contrast, there is no clear correlation between $B_{up}$ and $a$. We note that, at $t=0.15$, i.e. before tearing instability is triggered, most of the current sheets have lengths close to the wavelengths of initial fluctuations marked by the yellow shades. Moreover, panel (b1) shows that the upstream magnetic field is also quite consistent among all the current sheets, slightly above the initial fluctuation level ($\delta b\approx 0.17$). These results indicate that the generation of current sheets is dominated by the largest turbulence eddies in the system.
Figure 9: Probability distribution function (PDF) of $\cos(\theta_{ub})$ (a) and $\cos(\theta_{z^{\pm}})$ (b) at $t=0.3$. Here $\theta_{ub}$ is the angle between $\bm{u}$ and $\bm{b}$, and $\theta_{z^{\pm}}$ is the angle between $\bm{z^{+}}$ and $\bm{z^{-}}$. Background magnetic field was subtracted before calculating the angles. In each panel, blue bars correspond to regions outside current sheets and orange bars correspond to regions inside current sheets.
Figure 10: Scale-dependent alignment angles between (a) $\bm{u}$ and $\bm{b}$ and (b) $\bm{z^{+}}$ and $\bm{z^{-}}$. Solid curves with different colors correspond to different times. Dashed and dotted lines show $\propto l^{0.25}$ and $l^{0.1}$ for reference. (c) & (d): Same as (a) & (b) but the horizontal axis is $\lambda$ (perpendicular to local $\bm{\delta b}$) instead of $l$.
4 Discussion: Current sheets and scale-dependent dynamic alignment
[…]
In Figure 10, we show the scale-dependent alignment angles. Panels (a) and (b) show $\theta_{ub}(l)$ and $\theta_{z^{\pm}}(l)$, and panels (c) and (d) show $\theta_{ub}(\lambda)$ and $\theta_{z^{\pm}}(\lambda)$. Here $l=|\bm{l}|$ is non-directional spatial increment, and $\lambda$ is the spatial increment perpendicular to $\bm{\delta b}$, that is, the angle $\phi$ between $\bm{l}$ and $\delta\bm{b}(\bm{l})$ falls into the range $88^{\circ}-90^{\circ}$. The alignment angles are calculated using
$\theta_{ub}=\arcsin\left(\frac{\langle\left|\delta\bm{u}\times\delta\bm{b}\right|\rangle}{\langle\left|\delta\bm{u}\right|\left|\delta\bm{b}\right|\rangle}\right)$ (2)
and similar for $\theta_{z^{\pm}}$. We employ a Monte-Carlo method to evaluate the angles. That is, we randomly select two grid points $\bm{x_{1}}$ and $\bm{x_{2}}$ such that $\bm{l}=\bm{x_{2}}-\bm{x_{1}}$ and define $\delta\bm{b}=\bm{b}(\bm{x_{2}})-\bm{b}(\bm{x_{1}})$ and same for $\delta\bm{u}$ and $\delta\bm{z^{\pm}}$. We repeat this random process for $10^{8}$ times and calculate quantities like $\langle\left|\delta\bm{u}\times\delta\bm{b}\right|\rangle$, $\langle|\delta\bm{u}||\delta\bm{b}|\rangle$, and subsequently the alignment angles for the binned $l$ and $\phi$. Figure 10 confirm that both the alignment angles are smaller as we go toward smaller scales. In the classic SDDA models (S. Boldyrev, 2006; N. F. Loureiro & S. Boldyrev, 2017), the alignment angle scales as $\theta_{ub}\propto\lambda^{0.25}$, while in the later model by B. D. Chandran et al. (2015) the angles follow $\theta_{ub}\propto\lambda^{0.21}$ and $\theta_{z^{\pm}}\propto\lambda^{0.1}$. Although these previous results are not directly applicable to the 2D turbulence, we still plot two reference lines with slopes of 0.25 and 0.10 respectively in each panel of Figure 10 for visual assistance. Panels (a) and (b) show that, at early stage $t=0.15$, a single power law is observed for both the angles, and the slope for $\theta_{z^{\pm}}$ is steeper than that for $\theta_{ub}$. At later times, double power-law relations form for both angles with a break scale at $l\sim(1-2)\times 10^{-2}$. The large scale shows a slope close to $0.1$ and the small scale shows a slope close to $0.25$. Panels (c) and (d) also show a clear break at $\lambda\sim(1-2)\times 10^{-2}$. The power-law slopes for small scales are similar to the top panels, but the slopes for large scales are quite different from the top panels. What controls this break frequency and how the alignment angles change with scale is still unclear and requires future investigation. Another noteworthy point is that, for $\theta_{z^{\pm}}(l)$, the alignment angle first evolves to a low level at $t=0.15$ and then gradually rises. $\theta_{z^{\pm}}(\lambda)$ shows a similar evolution trend, though only for small scales. This is reminiscent of the evolution of Kurtosis($\bm{b}$) shown in Figure 4 and the residual energy shown in Figure 1 and it implies that $\theta_{z^{\pm}}$ is likely related to the evolution of turbulence current sheets.
Figure 11: (a) Second-order structure functions of magnetic field $S_{2}(\bm{b})$ as a function of spatial increment $l$ at $t=0.6$. Different curves correspond to different angles between $\delta\bm{b}(\bm{l})$ and $\bm{l}$. (b) Scale parallel ($\xi$) to $\delta\bm{b}$ as a function of scale perpendicular ($\lambda$) to $\delta\bm{b}$. (c) $\lambda/\xi$ as a function of $\lambda$.
Last, we quantify the anisotropy of the turbulence “eddies” using the method described in (N. Sioulas et al., 2024b). We apply the Monte-Carlo method to evaluate the second-order structure function of magnetic field while taking $\phi$ into consideration. As shown by Figure 11(a), $S_{2}(\bm{b})$ is clearly anisotropic between the directions parallel ($\xi$, $\phi=0$) and perpendicular ($\lambda$, $\phi=90^{\circ}$) to the local magnetic field increment. We numerically determine the function $\xi(\lambda)$ by equating $S_{2}(\bm{b},\xi)$ and $S_{2}(\bm{b},\lambda)$, and the result is shown in panel (b), which displays $\xi$ versus $\lambda$, and panel (c), which displays $\xi/\lambda$ versus $\lambda$. In general, $\xi>\lambda$, even at $t=0$, meaning that the turbulence eddies are stretched along the polarization direction. $\lambda/\xi$ is roughly scale-independent at $t=0$, with values 0.5-0.6, but becomes scale-dependent soon after the simulation starts. Similar to the Kurtosis shown in Figure 4 and $\theta_{z^{\pm}}$ shown in Figure 10, $\lambda/\xi$ drops to its lowest value at $t=0.15$ and then starts to rise, implying that it is related to the formation of turbulence current sheets. However, we emphasize that, a temporal correlation or concurrence does not ensure causal relation. The fact that current sheets, dynamic alignment, and turbulence eddy anisotropy evolve in a similar pattern does not mean that the current sheets are generated as a result of the dynamic alignment and eddy anisotropy. In fact, the estimated $\lambda/\xi$ (Figure 11) are much larger than the aspect ratio of the current sheets (Figure 8(a1-a3)). In addition, the filling factor of the current sheets is very low (<0.1<0.1, Figure 5) while the alignment angles and eddy anisotropy are evaluated over the whole simulation domain, meaning that the current sheet properties have negligible contribution in evaluation of the alignment angles and eddy anisotropy.
5 Summary
In this study, we conducted a high-resolution 2D simulation of balanced incompressible MHD turbulence. We comprehensively analyzed the properties of current sheets generated during the turbulence evolution. The major findings are summarized below:
  1. 1.
    The current sheets form much earlier ($t=0.15$) than one eddy turnover time ($t\approx 0.7$). The initial current sheet lengths are mostly comparable to the energy injection scales, i.e. they are controlled by the largest eddies in the system.
  2. 2.
    As the current sheets continue to thin, tearing instability onsets when the current sheet aspect ratio approaches $S^{-1/2}$.
  3. 3.
    Scale-dependent dynamic alignment is observed as shown by the alignment angles and eddy anisotropy. However, the alignment is much weaker than what is needed to explain the generated thin current sheets.
Our results indicate that there is no direct correspondence between the current sheet properties and SDDA. Therefore, it is necessary to revisit the reconnection-mediated turbulence model. Particularly, we need to find the proper way to relate the current sheet properties with the turbulence diagnostics.
CS is supported by NSF SHINE #2229566 and NASA ECIP #80NSSC23K1064. The numerical simulations were conducted on Derecho: HPE Cray EX System (https://doi.org/10.5065/qx9a-pg09) of Computational and Information Systems Laboratory (CISL), National Center for Atmospheric Research (NCAR) (allocation No. UCLA0063).

### v2

@@
 Finally, although the turbulence energy shows anisotropy between the directions parallel and perpendicular to the local magnetic field increment, we do not observe a direct correspondence between the shape of the current sheets and that of the turbulence “eddies.” These results suggest that one needs to be cautious when applying the scale-dependent dynamic alignment model to the analysis of current sheets in MHD turbulence.
-software: Matplotlib (J. D. Hunter, 2007), NumPy (C. R. Harris et al., 2020), Scikit-Learn (F. Pedregosa et al., 2011), LAPS (C. Shi et al., 2024)
-1 Introduction
-It has long been observed that the solar wind is highly turbulent, containing fluctuations on a vast span of scales (R.
-Bruno & V.
-Carbone, 2013).
+software: Matplotlib (Hunter2007Matplotlib), NumPy (harris2020array), Scikit-Learn (scikit-learn), LAPS (shi2024laps)
+I Introduction
+It has long been observed that the solar wind is highly turbulent, containing fluctuations on a vast span of scales (bruno2013solar).
 The solar wind turbulence is believed to be one of the major contributors to the heating and acceleration of solar wind and thus has been extensively investigated for decades (e.g.
-E.
-Marsch & C.-Y.
-Tu, 1990; B.
-Goldstein et al., 1995; R.
-Grappin & M.
-Velli, 1996; W.
-Matthaeus et al., 2005; C.
-H.
-Chen et al., 2012; O.
-Alexandrova et al., 2013; C.
-Shi et al., 2021; N.
-Sioulas et al., 2023).
+marsch1990spectral; goldstein1995properties; grappin1996waves; matthaeus2005spatial; chen2012three; alexandrova2013solar; shi2021alfvenic; sioulas2023magnetic).
 One of the key questions regarding solar wind turbulence is how its energy is dissipated.
 The widely-accepted picture is that, as the fluctuation energy cascades from large, MHD scales toward small, kinetic scales, various types of wave-particle interactions take effect, transferring energy from the turbulence to the particles (e.g.
-S.
-S.
-Cerri et al., 2017; J.
-He et al., 2019; M.
-M.
-Martinović et al., 2020; T.
-A.
-Bowen et al., 2022).
+cerri2017plasma; he2019direct; martinovic2020enhancement; bowen2022situ).
 In addition to these wave-particle interaction processes, intermittency naturally arises in turbulence systems, including neutral fluid (e.g.
-U.
-Frisch & G.
-Parisi, 1980) and plasma (e.g.
-L.
-Sorriso-Valvo et al., 2001).
+frisch1980fully) and plasma (e.g.
+sorriso2001intermittency).
 Intermittency refers to the sparse, irregular and strong fluctuations on top of the periodic and smooth part of the turbulence, and it breaks down the self-similarity of the turbulence.
-In MHD turbulence, intermittency mainly appears in the form of current sheets and vortices (M.
-Wan et al., 2016), and it is well-known that current sheets are usually accompanied by strong dissipation via magnetic reconnection (K.
-Osman et al., 2014).
+In MHD turbulence, intermittency mainly appears in the form of current sheets and vortices (wan2016intermittency), and it is well-known that current sheets are usually accompanied by strong dissipation via magnetic reconnection (osman2014magnetic).
 Therefore, intermittency may play a significant role in dissipation of MHD turbulence.
-In the solar wind, intermittency has been extensively observed (R.
-Bruno et al., 2001; W.
-H.
-Matthaeus et al., 2015; Z.
-Pei et al., 2016; H.
-Wu et al., 2023), and evidence shows that intermittency is co-located with regions of higher proton and electron temperatures in the solar wind (K.
-Osman et al., 2012; N.
-Sioulas et al., 2022b, a).
+In the solar wind, intermittency has been extensively observed (bruno2001identifying; matthaeus2015intermittency; pei2016influence; wu2023intermittency), and evidence shows that intermittency is co-located with regions of higher proton and electron temperatures in the solar wind (osman2012intermittency; sioulas2022statistical; sioulas2022preferential).
 The question then comes to how do we understand the generation of intermittency in MHD turbulence.
-In incompressible-MHD, phenomenological models were developed based on the idea of scale-dependent dynamic alignment (SDDA) (S.
-Boldyrev, 2005; J.
-Podesta et al., 2009; B.
-D.
-Chandran et al., 2015; A.
-Mallet et al., 2016; N.
-Sioulas et al., 2024a).
+In incompressible-MHD, phenomenological models were developed based on the idea of scale-dependent dynamic alignment (SDDA) (boldyrev2005spectrum; podesta2009scale; chandran2015intermittency; mallet2016measures).
 In this scenario, the small-scale turbulence “eddies” are assumed to be anisotropic three-dimensionally.
@@
 $\xi>\lambda$.
-A set of scaling relations between these lengths were derived, which were then coupled with the tearing instability theory with the assumption that the ratio $\xi/\lambda$ corresponds to the aspect ratio of the small-scale current sheets (S.
-Boldyrev, 2006; S.
-Boldyrev & N.
-F.
-Loureiro, 2017; N.
-F.
-Loureiro & S.
-Boldyrev, 2017; A.
-Mallet et al., 2017).
+A set of scaling relations between these lengths were derived, which were then coupled with the tearing instability theory with the assumption that the ratio $\xi/\lambda$ corresponds to the aspect ratio of the small-scale current sheets (boldyrev2006spectrum; boldyrev2017magnetohydrodynamic; loureiro2017role; mallet2017disruption).
 As the scale decreases, the aspect ratio may reach a certain critical value, after which the turbulence is dominated by the recursive collapse of current sheets, i.e.
-the so-called “reconnection-mediated” regime, associated with a steepened power spectrum (C.
-Dong et al., 2022).
+the so-called “reconnection-mediated” regime, associated with a steepened power spectrum (dong2022reconnection).
 Although both observations (e.g.
-J.
-Podesta et al., 2009; N.
-Sioulas et al., 2024a) and numerical simulations (e.g.
-J.
-Mason et al., 2006; J.
-Walker et al., 2018) have shown evidence of SDDA in plasma turbulence, there is a potential caveat in applying the SDDA theory to the reconnection in turbulence.
+podesta2009scale; sioulas2024scale) and numerical simulations (e.g.
+mason2006dynamic; walker2018influence) have shown evidence of SDDA in plasma turbulence, there is a potential caveat in applying the SDDA theory to the reconnection in turbulence.
 That is, the turbulence-generated current sheets are intermittent and thus are sparse in space instead of being space-filling like the turbulent eddies.
@@
 Nonetheless, we are still able to compute the aspect ratio of the 2D eddies $\xi/\lambda$ and verify whether they are related to the aspect ratio of the current sheets.
-The paper is organized as follows: In Section 2, we introduce the numerical methods and describe the simulation setup.
-In Section 3, we present the results, including the global evolution of turbulence properties (Section 3.1), the method to identify and analyze the intermittent current sheets (Section 3.2), a case study of an evolving current sheet (Section 3.3), and the statistics of the current sheet properties (Section 3.4).
-In Section 4, we discuss the results with a focus on comparing SDDA with current sheet properties.
-In Section 5, we summarize this work.
-2 Simulation setup
+The paper is organized as follows: In Section II, we introduce the numerical methods and describe the simulation setup.
+In Section III, we present the results, including the global evolution of turbulence properties (Section III.1), the method to identify and analyze the intermittent current sheets (Section III.2), a case study of an evolving current sheet (Section III.3), and the statistics of the current sheet properties (Section III.4).
+In Section IV, we discuss the results with a focus on comparing SDDA with current sheet properties.
+In Section V, we summarize this work.
+II Simulation setup
 We use the UCLA-Pseudo-Spectral (LAPS) code to conduct the simulation.
-LAPS is a Fourier-transform based pseudo-spectral MHD code (C.
-Shi et al., 2024).
+LAPS is a Fourier-transform based pseudo-spectral MHD code (shi2024laps).
 It has compressible and incompressible versions implemented in both 2D and 3D.
@@
 Thus the (magnetic) Reynolds number is estimated to be $S_{0}=\lambda\delta b/\eta\approx 4\times 10^{4}$.
-We conducted tests with different amplitudes of dissipation and the values $\eta=\nu=5\times 10^{-7}$ are chosen such that the dissipation is strong enough to effectively suppress the numerical error induced by Gibbs phenomenon while not too strong so that a significant number of tearing-unstable current sheets are generated (M.
-Wan et al., 2013).
+We conducted tests with different amplitudes of dissipation and the values $\eta=\nu=5\times 10^{-7}$ are chosen such that the dissipation is strong enough to effectively suppress the numerical error induced by Gibbs phenomenon while not too strong so that a significant number of tearing-unstable current sheets are generated (wan2013generation).
 The eddy-turnover time $\tau_{nl}=\lambda/z_{\lambda}\approx 0.7$.
 Thus, we run the simulation to $t=1.0$, roughly 1.4 eddy-turnover times.
-Figure 3: Evolution of $J_{z}$ in the subdomain $x\in[0.2,0.7]$, $y\in[0,0.5]$. In Panel (c), the green box marks the current sheet analyzed in detail in Section 3.3 and Figure 6.
-3 Results
-3.1 Evolution of turbulence properties
+Figure 3: Evolution of $J_{z}$ in the subdomain $x\in[0.2,0.7]$, $y\in[0,0.5]$. In Panel (c), the green box marks the current sheet analyzed in detail in Section III.3 and Figure 6.
+III Results
+III.1 Evolution of turbulence properties
 In Figure 1, we show the time evolution of various quantities.
 Panel (a) shows that $\sigma_{c}$ (blue) remains roughly zero throughout the simulation, while $\sigma_{r}$ (orange) decreases rapidly at the early stage ($t\leq 0.15$), then increases a little bit and finally starts to decay slowly.
-The initial rapid decrease of $\sigma_{r}$ was observed in 3D simulations with similar setup (C.
-Shi et al., 2025).
+The initial rapid decrease of $\sigma_{r}$ was observed in 3D simulations with similar setup (shi2025evolution).
 In Panel (b), we show the evolution of $E_{v}$ (blue), $E_{b}$ (orange), $E_{+}$ (green), and $E_{-}$ (red) respectively.
@@
 This has been confirmed by a large number of numerical simulations (e.g.
-W.-C.
-Müller & R.
-Grappin, 2005; S.
-Boldyrev et al., 2011; C.
-Shi et al., 2025) and observations (e.g.
-J.
-Podesta et al., 2007; C.
-Chen et al., 2013; C.
-Shi et al., 2021).
+muller2005spectral; boldyrev2011spectral; papini2019can; shi2025evolution) and observations (e.g.
+podesta2007spectral; chen2013residual; shi2021alfvenic).
 The slopes for $\bm{z^{\pm}}$ are similar and take values between 3/2-3/2 and 5/3-5/3.
@@
 Tearing instability has been triggered in some of the current sheets, indicated by the generation of plasmoids.
-One tearing-unstable current sheet is marked by the green box and this current sheet will be analyzed in detail in Section 3.3.
+One tearing-unstable current sheet is marked by the green box and this current sheet will be analyzed in detail in Section III.3.
 Panel (d) shows $t=0.6$, corresponding to the vertical dashed line in Figure 1, when the turbulence is fully developed.
@@
 Interestingly, it is observed that the Kurtosis reaches maximum at $t=0.15$ and then gradually decay.
-Combined with Figure 1, it indicates that the increase of Kurtosis is associated with the current sheet formation process, and the later reconnection decreases the Kurtosis.
-3.2 Identification of current sheets
+Combined with Figure 1, it indicates that the increase of Kurtosis is associated with the current sheet formation process, and the later reconnection results in a decrease of the Kurtosis.
+III.2 Identification of current sheets
 To quantify the properties of the current sheets, it is necessary to accurately identify the current sheets.
-We develop an algorithm similar to that described in (V.
-Zhdankin et al., 2013), and we briefly describe our algorithm here.
+We develop an algorithm similar to that described in (zhdankin2013statistical), and we briefly describe our algorithm here.
 We start from finding the grid point corresponding to $\max(\left|J_{z}\right|)$ throughout the whole domain, excluding regions already marked as current sheets.
@@
 It shows that the algorithm works well, and most of the identified current sheets exist between the interacting large-scale “eddies”, i.e.
-near the eddy boundaries.
+near the eddy boundaries, as pointed out by papini2019can in their Hall-MHD and hybrid Particle-in-Cell simulations.
 However, we note that, as the turbulence develops, there are many regions with strong current density but are “broken” instead of being coherent structures.
@@
 In each panel, top plot shows $J_{z}$ and bottom plot shows $\omega_{z}$.
-In panels (a) & (b), the green shade is the automatically identified current sheet region using the algorithm described in Section 3.2.
-The white and red lines correspond to the segments of the current sheet given by the recursive PCA algorithm described in Section 3.3.
-3.3 Case study of a current sheet
+In panels (a) & (b), the green shade is the automatically identified current sheet region using the algorithm described in Section III.2.
+The white and red lines correspond to the segments of the current sheet given by the recursive PCA algorithm described in Section III.3.
+III.3 Case study of a current sheet
 We aim to accurately quantify the geometry of these current sheets, including their lengths and thicknesses.
-Therefore, a principal component analysis (PCA) based algorithm is developed, with the aid of PCA functions implemented in the Python package Scikit-Learn (F.
-Pedregosa et al., 2011).
+Therefore, a principal component analysis (PCA) based algorithm is developed, with the aid of PCA functions implemented in the Python package Scikit-Learn (scikit-learn).
 For a set of spatial coordinates $(x_{i},y_{i})$ ($i=0,\cdots N-1$), PCA determines two orthogonal unit vectors, along which the set of coordinates has maximum and minimum variances respectively.
@@
 We note that, the Lundquist number $S_{L}=6.5\times 10^{4}$ of this current sheet is not significantly larger than the critical Lundquist number, approximately $O(10^{4})$, for the Sweet-Parker type current sheet to be tearing-unstable (e.g.
-D.
-Biskamp, 1986; N.
-Loureiro et al., 2007; C.
-Shi et al., 2018).
-In the nearly-collisionless space plasma with much larger Lundquist numbers, the growth of linear tearing mode may happen earlier, when the aspect ratio $a/L$ approaches $S_{L}^{-1/3}$ instead of $S_{L}^{-1/3}$ (F.
-Pucci & M.
-Velli, 2013; A.
-Tenerani et al., 2015, 2016).
-3.4 Statistics of current sheets
+biskamp1986magnetic; loureiro2007instability; shi2018marginal).
+In the nearly-collisionless space plasma with much larger Lundquist numbers, the growth of linear tearing mode may happen earlier, when the aspect ratio $a/L$ approaches $S_{L}^{-1/3}$ instead of $S_{L}^{-1/2}$ (pucci2013reconnection; tenerani2015magnetic; tenerani2016ideally; del2016ideal).
+III.4 Statistics of current sheets
 Figure 8: Statistics of the current sheet properties.
@@
 In the right column, the yellow shades mark $1/16\leq L\leq 1/8$.
-We apply the analysis described in Section 3.3 to all the identified current sheets and present the results in Figure 8.
+We apply the analysis described in Section III.3 to all the identified current sheets and present the results in Figure 8.
 Top to bottom rows show time moments $t=0.15$, $0.3$, and $0.6$ respectively.
@@
 We can see that for all the three moments, the data points show a negative correlation between $a/L$ and $S_{L}$, and they roughly obey $a/L\sim S_{L}^{-1/2}$.
-In the middle and right columns, we plot reference lines $B_{up}\propto a^{1/4}$ and $L\propto a^{3/4}$, which are predictions by the SDDA theory (S.
-Boldyrev, 2006).
-There is a strong positive correlation between $L$ and $a$, and they may roughly follow the power-law relation $L\propto a^{3/4}$.
-In contrast, there is no clear correlation between $B_{up}$ and $a$.
-We note that, at $t=0.15$, i.e.
+In the middle column, we plot reference lines $B_{up}\propto a^{1/4}$, which is the prediction by the SDDA theory (boldyrev2006spectrum).
+One can see that there is no clear correlation between $B_{up}$ and $a$.
+In contrast, the right column shows a strong positive correlation between $L$ and $a$.
+For reference, we plot $L\propto a^{2}$, which is the scaling relation given by Sweet-Parker model, i.e.
+$a/L\sim(LB_{up}/\eta)^{-1/2}$, assuming that the upstream Alfvén speed is independent of $a$.
+Similarly, for ideal-tearing scaling relation $a/L\sim(LB_{up}/\eta)^{-1/3}$, we will get $L\propto a^{3/2}$.
+Note that, SDDA theory predicts $L\propto a^{3/4}$, much shallower than what is shown here.
+We point out that, at $t=0.15$, i.e.
 before tearing instability is triggered, most of the current sheets have lengths close to the wavelengths of initial fluctuations marked by the yellow shades.
 Moreover, panel (b1) shows that the upstream magnetic field is also quite consistent among all the current sheets, slightly above the initial fluctuation level ($\delta b\approx 0.17$).
-These results indicate that the generation of current sheets is dominated by the largest turbulence eddies in the system.
+These results indicate that the generation of current sheets is dominated by the largest turbulence eddies in the system (papini2019can; khan2025does), and their subsequent evolution is dominated by the recursive tearing instability (shibata2001plasmoid; tenerani2015magnetic).
 Figure 9: Probability distribution function (PDF) of $\cos(\theta_{ub})$ (a) and $\cos(\theta_{z^{\pm}})$ (b) at $t=0.3$.
@@
 (c) & (d): Same as (a) & (b) but the horizontal axis is $\lambda$ (perpendicular to local $\bm{\delta b}$) instead of $l$.
-4 Discussion: Current sheets and scale-dependent dynamic alignment
+IV Discussion: Current sheets and scale-dependent dynamic alignment
 […]
@@
 Figure 10 confirm that both the alignment angles are smaller as we go toward smaller scales.
-In the classic SDDA models (S.
-Boldyrev, 2006; N.
-F.
-Loureiro & S.
-Boldyrev, 2017), the alignment angle scales as $\theta_{ub}\propto\lambda^{0.25}$, while in the later model by B.
-D.
-Chandran et al.
-(2015) the angles follow $\theta_{ub}\propto\lambda^{0.21}$ and $\theta_{z^{\pm}}\propto\lambda^{0.1}$.
+In the classic SDDA models (boldyrev2006spectrum; loureiro2017role), the alignment angle scales as $\theta_{ub}\propto\lambda^{0.25}$, while in the later model by chandran2015intermittency the angles follow $\theta_{ub}\propto\lambda^{0.21}$ and $\theta_{z^{\pm}}\propto\lambda^{0.1}$.
 Although these previous results are not directly applicable to the 2D turbulence, we still plot two reference lines with slopes of 0.25 and 0.10 respectively in each panel of Figure 10 for visual assistance.
@@
 $\theta_{z^{\pm}}(\lambda)$ shows a similar evolution trend, though only for small scales.
-This is reminiscent of the evolution of Kurtosis($\bm{b}$) shown in Figure 4 and the residual energy shown in Figure 1 and it implies that $\theta_{z^{\pm}}$ is likely related to the evolution of turbulence current sheets.
+This is reminiscent of the evolution of Kurtosis($\bm{b}$) shown in Figure 4 and the residual energy shown in Figure 1 and it implies that $\theta_{z^{\pm}}$ is likely affected by the evolution of turbulence current sheets.
 Figure 11: (a) Second-order structure functions of magnetic field $S_{2}(\bm{b})$ as a function of spatial increment $l$ at $t=0.6$.
@@
 (c) $\lambda/\xi$ as a function of $\lambda$.
-Last, we quantify the anisotropy of the turbulence “eddies” using the method described in (N.
-Sioulas et al., 2024b).
+Last, we quantify the anisotropy of the turbulence “eddies” using the method described in (sioulas2024higher).
 We apply the Monte-Carlo method to evaluate the second-order structure function of magnetic field while taking $\phi$ into consideration.
@@
 $\lambda/\xi$ is roughly scale-independent at $t=0$, with values 0.5-0.6, but becomes scale-dependent soon after the simulation starts.
-Similar to the Kurtosis shown in Figure 4 and $\theta_{z^{\pm}}$ shown in Figure 10, $\lambda/\xi$ drops to its lowest value at $t=0.15$ and then starts to rise, implying that it is related to the formation of turbulence current sheets.
+Similar to the Kurtosis shown in Figure 4 and $\theta_{z^{\pm}}$ shown in Figure 10, $\lambda/\xi$ drops to its lowest value at $t=0.15$ and then starts to rise, implying that it is more or less related to the formation of turbulence current sheets.
 However, we emphasize that, a temporal correlation or concurrence does not ensure causal relation.
 The fact that current sheets, dynamic alignment, and turbulence eddy anisotropy evolve in a similar pattern does not mean that the current sheets are generated as a result of the dynamic alignment and eddy anisotropy.
-In fact, the estimated $\lambda/\xi$ (Figure 11) are much larger than the aspect ratio of the current sheets (Figure 8(a1-a3)).
-In addition, the filling factor of the current sheets is very low (<0.1<0.1, Figure 5) while the alignment angles and eddy anisotropy are evaluated over the whole simulation domain, meaning that the current sheet properties have negligible contribution in evaluation of the alignment angles and eddy anisotropy.
-5 Summary
+In fact, the estimated $\lambda/\xi$ (Figure 11) are much larger than the aspect ratio of the current sheets (Figure 8(a1-a3)), which is mostly less than 0.1.
+In addition, the filling factor of the current sheets is very low (<0.1<0.1, Figure 5) while the alignment angles and eddy anisotropy are evaluated over the whole simulation domain, meaning that the current sheet properties should have negligible contribution in the evaluation of the alignment angles and eddy anisotropy.
+V Summary
 In this study, we conducted a high-resolution 2D simulation of balanced incompressible MHD turbulence.
@@
     As the current sheets continue to thin, tearing instability onsets when the current sheet aspect ratio approaches $S^{-1/2}$.
+This process happens in a recursive way and generates smaller and smaller current sheets.
   3. 3.
@@
 Particularly, we need to find the proper way to relate the current sheet properties with the turbulence diagnostics.
-CS is supported by NSF SHINE #2229566 and NASA ECIP #80NSSC23K1064.
+CS is supported by NSF SHINE #2548299 and NASA ECIP #80NSSC23K1064.
 The numerical simulations were conducted on Derecho: HPE Cray EX System (https://doi.org/10.5065/qx9a-pg09) of Computational and Information Systems Laboratory (CISL), National Center for Atmospheric Research (NCAR) (allocation No.